OpenStudy (anonymous):

does 5/8 and 6/12 form a proportion

OpenStudy (anonymous):

no

OpenStudy (anonymous):

To find out you to cross multiplication. so...\[8*12=96 \]

OpenStudy (anonymous):

\[5*6=30\]

OpenStudy (anonymous):

Since they arent equal then it means it isnt proportionate.

OpenStudy (anonymous):

thanks

hero (hero):

If it forms a proportion, then after cross multiplication, both sides would be equal. For example, if you had \[\frac{3}{8} = \frac{9}{24}\] After cross multiplication you get (3)(24) = (9)(8) 72 = 72

hero (hero):

Let try 5/8 and 6/12 to see if it forms a proportion: \[\frac{5}{8} = \frac{6}{12}\] After cross multiplication: (5)(12) = (8)(6) 60 = 48? False

hero (hero):

Also we know it is not a proportion because 6/12 = 50% 5/8 > 50% The percentages of both fractions must be equal in order for a proportion to exist.